# KCL-PH-TH/2026-03

### v2

@@ IV Axion-like fields from the $\mathcal{H}$-Torsion
and integrating by parts in the last (the $\lambda$-dependent ) term on the right-hand of (102), one can show that this higher-order term can be written in the form:
$\int\nabla_{\mu}b\nabla_{\nu}bR^{\mu\nu}\sqrt{-g}\,d^{4}x=\underbrace{\int\nabla_{\mu}(\nabla_{\nu}b\nabla^{\mu}\nabla^{\nu}b-\nabla^{\mu}b\Box b)\sqrt{-g}\,d^{4}x}_{=\rm Boundary\ term}+\int\left[(\Box b)^{2}-(\nabla_{\mu}\nabla_{\nu}b)(\nabla^{\mu}\nabla^{\nu}b)\right]\sqrt{-g}\,d^{4}x,$ (104)
where $\Box=\nabla_{\rho}\nabla^{\rho}$ is the (torsion-free) gravitationally covariant d’ Alembertian, and the boundary terms are dropped, assuming standard boundary conditions for fields in this Universe.
Variation of the effective action (102) with respect to the metric gives equations of the form:
@@
where $G_{\mu\nu}$ is the Einstein tensor, ${T_{\rm kin}^{b}}_{\mu\nu}$ is the axion stress-energy tensor (corresponding to the axion kinetic term only in (102)),
@@
$\mathcal{C_{\mu\nu}}$ is (a variant of) the Cotton tensor (or more precisely the C-tensor) [25, 26, 23], given as
@@
and $\Theta_{\mu\nu}$ is the tensor that arises after variation of $\partial_{\mu}b\partial_{\nu}bR^{\mu\nu}$, given as
$\Theta_{\mu\nu}=\nabla_{\sigma}\big((\nabla_{\mu}\nabla_{\nu}b)\nabla^{\sigma}b\big)-2R_{\sigma(\mu}(\nabla_{\nu)}b)(\nabla^{\sigma}b)-g_{\mu\nu}\Big[(\nabla^{\sigma}b)\nabla_{\sigma}\Box b+\frac{1}{2}(\Box b)^{2}+\frac{1}{2}(\nabla_{\sigma}\nabla_{\rho}b)(\nabla^{\sigma}\nabla^{\rho}b)\Big].$ (108)
The parameter $q$ in (IV) plays the rôle of the reduced Planck constant, renormalized by the $\alpha^{\prime}$ corrections. So from now on, we replace (in all the subsequent analysis)
@@
Variation with respect to $b$ gives the scalar equation of motion:
@@
with $\mathcal{K}^{\mu}$ being the (gravitational) Chern-Simons current,
@@
We next proceed to discuss the effects of the higher-order term (104) on the running-vacuum GW-condensate inflationary scenario of [2, 7].
V Stringy-Running-Vacuum-Model Inflation
Having arrived at the effective action (102) to fourth-derivative order, we are now well equipped to revisit the stringy running vacuum model cosmology (StRVM) [1, 2, 3, 4, 5], and see how the higher-order terms affect the relevant conclusions on the physics of the associated inflation. To this end we should first demonstrate that inflation is an admissible solution of the pertinent cosmological equations of motion, as was the case of the initial StRVM. The scalar equation of motion (110) takes the form of a conserved current equation:
@@ V Stringy-Running-Vacuum-Model Inflation
On assuming a (spatially-flat) background of an expanding isotropic and homogeneous Friedman-Lemaitre-Robertson-Walker (FLRW) universe, with scale factor $a(t)$, in which the axion-like fields depend only on cosmic time $t$, $b=b(t)$, the expression for the current simplifies to
@@
because $\partial_{0}b=\dot{b}$ and $\partial^{0}b=-\dot{b}$ (in our metric signature convention). So we have that
@@
The constant $C$ can be set to 0 if one considers an inflationary spacetime [2]. Furthermore, on an FLRW background, we have that during inflation, $a\sim\exp(H\,t)$, with $H=H_{I}\simeq{\rm constant}$ the approximately constant Hubble parameter, for which the cosmological data imply [42]:
@@
Thus, the time-time component of the Ricci tensor equals approximately:
@@
In the presence of primordial GW-perturbations the gravitational CS (gCS) anomaly term in (102) can condense at the end of the axion-dominated (pre-inflationary) stiff era  [2]. This can drive the Universe to an inflationary phase, characterised by approximately constant gCS-anomaly condensate and Hubble parameter $H$, and a linear axion $b$ potential in (102), breaking the axion shift symmetry. This leads to a metastable inflation [2, 7, 8], due to the existence of imaginary parts in the condensate [8], which lead to a finite-life time inflationary era, and eventual exit from it. However, although, from a purely dynamical systems point of view [7, 43] it appears that it is the $b$-field that drives inflation, this is deceptive. The condensate itself is a non-linear function of the (slowly varying) Hubble parameter $H$, depending on fourth- (and higher (even)-) order powers of it. In this sense, it is such non-linearities of the gravitational sector that drive an inflation of RVM type [15, 14, 13, 10, 9]. [^11] For completeness, we mention that the inflationary exit in such cosmologies, which in general do not require external inflaton fields, is characterised by prolonged reheating periods [44, 45], and perhaps early-matter dominance epoch, preceeding standard radiation. This may affect primordial-black-hole populations and thus gravitational-wave profiles during early radiation era, with potentially detectable signatures in future intrerferometers [46, 47].
To estimate the gCS condensation, the authors of [7], whose approach we follow here, considered tensor perturbations in the context of the effective gravitational Lagrangian (102), with $\lambda=0$. They quantize them within an effective weak gravitational field theory action [48, 49], following a canonical-quantization approach to estimate the condensate, to leading (linear) order in a perturbative expansion in the small parameter $\dot{b}H^{2}\ll 1\,.$
In the presence of the $\lambda$-dependent correction terms (104) in (102), higher-order corrections proportional to $(\dot{b})^{2}$ are expected in the condensate, which will be shown below to be subleading. For the $\lambda=0$ case, as argued in [2], and also confirmed by the dynamical-system analysis of the StRVM inflation of [7], and will also be reviewed below, when there is a gravitational anomaly condensate, a spontaneous-Lorentz-violating solution of the $b$-axion equations of motion exists in which $\langle\mathcal{K}^{0}\rangle\simeq{\rm constant}$. We parametrize such constant quantity by  [2, 7]
@@
during inflation, where $\varepsilon$ (not to be confused with $\epsilon$ in (IV)) is a numerical coefficient, which in the study of [8] is found to be of order $\mathcal{O}(10^{-2}$), and the Hubble parameter $H\approx H_{I}=\rm constant$, satisfying (120). As we shall demonstrate below, the order of magnitude of (123) will not be affected by the presence of the $\lambda$-dependent term in (102), whose contributions to gCS condensate are extremely suppressed.
To see this we should use perturbation theory in $\dot{b}$ when evaluating the gCS condensate, following [7]. To this end, formally we consider weak tensor (GW) perturbations to the FLRW metric, and work in the transverse-traceless (TT) gauge [48, 49], which implies that only their spatial components, $h_{ij}$, $i,j=1,2,3$, are present in the (gauge-fixed) effective action. The respective (3+1)-dimensional line-element is of the form:
@@
with $a(t)$ the scale factor.
In the helicity ($L,R$) basis, the metric reads:
$g_{\mu\nu}=\begin{pmatrix}-1&0&0&0\\ 0&a^{2}(t)\left(1+\frac{h_{L}(t,z)+h_{R}(t,z)}{\sqrt{2}}\right)&ia^{2}(t)\left(1+\frac{h_{L}(t,z)-h_{R}(t,z)}{\sqrt{2}}\right)&0\\ 0&ia^{2}(t)\left(1+\frac{h_{L}(t,z)-h_{R}(t,z)}{\sqrt{2}}\right)&a^{2}(t)\left(1-\frac{h_{L}(t,z)+h_{R}(t,z)}{\sqrt{2}}\right)&0\\ 0&0&0&a^{2}(t)\end{pmatrix}$ (125)
@@
We consider the GW perturbations $h_{ij}$ to be small, propagating along the $z$ direction for concreteness. We thus obtain the following expression for the gCS condensate, in the linearised approximation, in conformal time (the derivative with respect to which are denoted by primes) [50, 51, 7]:
@@
As discussed in [7], the quantity inside the square brackets $\left[\dots\right]$ on the right-hand-side is independent of the scale factor $a$ of the Universe, thus the gravitational CS condensate (127) is inversely proportional to the fourth power of the scale factor $a$.
To estimate (127), we first need to determine solutions to the linearized GW equations, in the modified effective action (102), which read:
@@
We observe from (V), (V) that the presence of the Chern-Simons coupling causes the equations for left and right waves to be different, a phenomenon known as cosmological birefringence. On the other hand, as expected, the $\lambda$-dependent corrections (104) do not exhibit that property. Thus, such terms alone (i.e. in the absence of the gravitational anomaly CS term, when $\epsilon=0$) yield zero contributions to (127). Of course, as we shall argue below, there are non-trivial, but suppressed $\lambda$-dependent corrections if $\epsilon\neq 0$.
To estimate those corrections we should first solve, following [7, 8] the above equations, in order to determine the so-called mode functions for the two GW polarizations. Then, upon canonical quantization of $h_{L,R}$, which now become quantum operators $\widehat{h}_{L,R}$, we can estimate the vacuum expectation value of (127) with respect to an appropriate quantum vacuum state, which, as in the $\lambda=0$ case of [7, 8], is taken to be the Bunch-Davies vacuum. This procedure will lead to a non-vanishing vacuum, approximately constant, expectation value for the Pontryagin term (127) during the inflationary period (indicated by the suffix “II”) [2, 7, 8]:
@@
which, due to the above discussion, is expected to be proportional to $\epsilon$.
Below we provide a rather heuristic approach for the estimation of the order of the $\lambda$-dependent-term contribution to the gCS condensate (for a more rigorous approach, see Appendix D). To this end, we first assume an approximately constant $\dot{b}$ during inflation, in which $H\simeq H_{I}={\rm constant}$, with the magnitude of $H_{I}$ given by observations [42], (120). In such a case, terms involving $\ddot{b}$ in the chiral GW equations of motion (V), (V) can be ignored. By redefining
@@
the equations (V), (V), after straightforward manipulations, can be approximated by:
@@
As discussed in Appendix D (cf. (188)), for the StRVM Cosmological model of [2, 4, 7, 8], the quantity $\lambda\kappa^{2}\dot{b}^{2}=\mathcal{O}(10^{-11})\ll 1$. On the other hand, upon taking the Fourier transform of the GW linear polarizations $h_{L,R}$, assuming propagation along the $z$ direction, we may replace the $\frac{1}{a}\partial_{z}$ by $i\frac{k}{a}$, where $k$ is the magnitude of the momentum vector of the GW polarization. We then have [2, 7] $\frac{k}{a}\lesssim\mu$, where $\mu$ is the UV cutoff, identified with the string scale [19] $M_{s}$, in the context of the StRVM, (100). For phenomenological reasons, specifically in order for the inflation in the StRVM to have a lifetime in the order of 50-60 e-foldings [42], the following constraint must be in operation, as implied by a dynamical-system analysis of the StRVM condensate-induced inflation [7],
@@
which is also in agreement with the study in [2].
On account of (100), (134), (123) and (120), then, we obtain that the dimensionless terms $\epsilon(\frac{k}{a})\kappa^{2}\dot{b}\lesssim 8.3\times 10^{-7}$. Thus the $\lambda$-dependent terms on the right-hand side of the wave equations (V), (V) may be ignored. In this approximation, the resulting equations, which are expressed in terms of $a^{\lambda}(t)$ ((131)), become equivalent to the corresponding wave equations of [7]. Following, then, that analysis, we may estimate the corresponding parity-violating condensate (127) in our modified gravitational theory (102), as:
@@
where we kept first-order terms in a series expansion in powers of $\lambda$, and we took into account the aforementioned property of the gravitational CS anomaly condensate to depend on the inverse fourth power of the (rescaled) $a^{\lambda}$ ((131)). From (135), we therefore observe that the $\lambda$-dependent corrections will be suppressed compared to the $\lambda=0$ results of [2, 7, 8], and thus the corresponding results, regarding the inflationary era, remain valid to an excellent approximation.
We now remark that in the $\lambda=0$ case, which constitutes the zeroth order correction in the $\lambda$ (i.e. Regge-slope $\alpha^{\prime}$ expansion in the string effective action [19]), the argumentation of [6], which was confirmed by the detailed analysis of [8], indicates that the gCS condensate (130) is proportional to the proper number density of GW sources $\mathcal{N}_{I}$ during the RVM inflationary era. In [7, 8] the Chern-Simons condensate has been estimated within a weak quantum-gravity path-integral formalism about a FLRW background.
In view of our result (135), then, in our modified effective action (101), the corresponding gCS anomaly condensate will be given by:
@@
to leading order in the slow-roll parameter $\alpha^{\prime}\,\kappa\,H\,{\dot{\overline{b}}}\,\ll\,1$ of the double perturbative expansion, in powers of $\alpha^{\prime}$ and $\dot{b}$, where $H$ is the approximately constant Hubble parameter during inflation, satisfying the observational bound (120).
As already mentioned the validity of (109) is understood. Thus, from (136), one observes that the effects of the presence of the $\lambda$-dependent terms in the modified gravitational action (102), as compared to the $\lambda=0$ case of [2, 4, 7], is merely a very mild screening (reduction) of the strength of the effect of the GW source terms $\mathcal{N}_{I}^{\prime}\simeq\mathcal{N}_{I}(1-2\lambda\,\kappa^{2}\,\dot{b}^{2})$.
The quantity $\mu$ in (136) is an UV cutoff of the GW modes, and the $\dots$ denote corrections implied by the last term of the integrand of (102), which we have just seen that they are subleading.
The existence of this condensate means that, at a quantum level, the gravitational Chern-Simons term in the effective action can be expanded about the condensate as
@@
where the $\int d^{4}x\sqrt{-g}:b\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}:$ represent quantum fluctuations, for which $\langle:\dots:\rangle=0$. As explained in [7], we can then write a “re-classicalised” effective action that has the form
$S_{E}=\int\bigg[\frac{1}{2q^{2}}R-\frac{1}{2}\partial_{\mu}b\partial^{\mu}b\,+\,\frac{\epsilon}{8}b\langle R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\rangle+\frac{\epsilon}{8}:b\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}:-\frac{\lambda}{2}\partial_{\mu}b\partial_{\nu}bR^{\mu\nu}\bigg]\sqrt{-g}\,d^{4}x.$ (138)
@@
$\frac{1}{2q^{2}}G_{\mu\nu}+\frac{\epsilon}{16}\langle R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\rangle g_{\mu\nu}=\frac{\epsilon}{2}\mathcal{C}_{\mu\nu}+\frac{\lambda}{2}\Theta_{\mu\nu}+\frac{1}{2}{T_{\rm kin}^{b}}_{\mu\nu},$ (139)
while the scalar equation of motion read:
@@
On expanding the effective action (137) about the condensate, implies in general
@@
where, as already mentioned, the $\dots:$ denote quantum fluctuations, characterised by a zero vacuum expectation value, $\langle:\dots:\rangle=0$. [^12] In the case of the gravitational CS term, the quantum fluctuations do not contribute to the equations of motion of graviton in the FLRW background.
A constant condensate $\mathcal{K}^{0}$ (which spontaneously violates Lorentz symmetry [2], if we view the condensate as a vacuum expectation value of the topological current $\langle\mathcal{K}^{0}\rangle=\rm constant\neq 0$) is consistent with inflation, characterized by $H=\dot{a}/a={\rm constant}$ and $\ddot{a}/a={\rm constant}$, which, on account of (122), would also imply $\dot{b}={\rm constant}$, as in the case of [2]. The reader should notice that (123) implies a $H^{4}$ scaling of the Chern-Simons condensate (136) during inflation.
On concentrating on the condensate (which behaves classically), and ignoring the fluctuations, we obtain from (113), (112):

### v3

@@ IV Axion-Like Fields from the $\mathcal{H}$-Torsion
and integrating by parts in the last (the $\lambda$-dependent ) term on the right-hand of (102), one can show that this higher-order term can be written in the following form:
$\int\nabla_{\mu}b\nabla_{\nu}bR^{\mu\nu}\sqrt{-g}\,d^{4}x$ $=\underbrace{\int\nabla_{\mu}(\nabla_{\nu}b\nabla^{\mu}\nabla^{\nu}b-\nabla^{\mu}b\Box b)\sqrt{-g}\,d^{4}x}_{=\rm Boundary\ term}+$
$+\int\left[(\Box b)^{2}-(\nabla_{\mu}\nabla_{\nu}b)(\nabla^{\mu}\nabla^{\nu}b)\right]\sqrt{-g}\,d^{4}x,$ (104)
where $\Box=\nabla_{\rho}\nabla^{\rho}$ is the (torsion-free) gravitational covariant of d’Alembertian, and the boundary terms are dropped, assuming standard boundary conditions for fields in this Universe.
It is also important to notice that as a result of this $\lambda$-dependent term, the kinetic terms of the axion field $b(x)$ in the effective action (102) appear to correspond to a $\mathcal{O}(\alpha^{\prime})$ “corrected metric tensor”
[…]
of the type appearing in the seminal papers of Friedan [42, 43], which constitutes the basis for the conformal $\sigma$-model approach to target-space string effective actions [29] and is perturbatively equivalent to the $S$-matrix approach [28].
Variation in the effective action (102) with respect to the metric gives equations of the following form:
@@
where $G_{\mu\nu}$ is the Einstein tensor, ${T_{\rm kin}^{b}}_{\mu\nu}$ is the axion stress energy tensor (corresponding to the axion kinetic term only in (102))
@@
$\mathcal{C_{\mu\nu}}$ is a variant of the Cotton tensor (or more precisely the C-tensor) [25, 26, 23], given as
@@
and $\Theta_{\mu\nu}$ is the tensor that arises after variation in $\partial_{\mu}b\partial_{\nu}bR^{\mu\nu}$, given as
$\Theta_{\mu\nu}$ $=\nabla_{\sigma}\big((\nabla_{\mu}\nabla_{\nu}b)\nabla^{\sigma}b\big)-2R_{\sigma(\mu}(\nabla_{\nu)}b)(\nabla^{\sigma}b)$-
$-g_{\mu\nu}\Big[(\nabla^{\sigma}b)\nabla_{\sigma}\Box b+\frac{1}{2}(\Box b)^{2}+\frac{1}{2}(\nabla_{\sigma}\nabla_{\rho}b)(\nabla^{\sigma}\nabla^{\rho}b)\Big].$ (109)
The parameter $q$ in (IV) plays the role of the reduced Planck constant, renormalized by the $\alpha^{\prime}$ corrections. So, from now on, we replace the following in all subsequent analyses:
@@
Variation with respect to $b$ gives the following scalar equation of motion:
@@
with $\mathcal{K}^{\mu}$ being the gravitational Chern–Simons current, as follows:
@@
We next proceed to discuss the effects of the higher-order term (IV) on the running-vacuum GW-condensate inflationary scenario of [2, 7].
V Stringy Running Vacuum Model Inflation
Having arrived at the effective action (102) to the fourth-derivative order, we are now well equipped to revisit the stringy running vacuum model cosmology (StRVM) [1, 2, 3, 4, 5] and see how the higher-order terms affect the relevant conclusions on the physics of the associated inflation. To this end, we shall first demonstrate that inflation is an admissible solution of the pertinent cosmological equations of motion, as was the case for the initial StRVM. The scalar equation of motion (111) takes the form of a conserved current equation as follows:
@@ V Stringy Running Vacuum Model Inflation
Upon assuming a spatially flat background of an expanding isotropic and homogeneous Friedman–Lemaitre–Robertson–Walker (FLRW) Universe, with scale factor $a(t)$, in which the axion-like fields depend only on cosmic time $t$, $b=b(t)$, the expression for the current simplifies to
@@
because $\partial_{0}b=\dot{b}$ and $\partial^{0}b=-\dot{b}$ in our metric signature convention. So, we have
@@
The constant $C$ can be set to 0 if one considers an inflationary spacetime [2]. Furthermore, on an FLRW background, we have that during inflation $a\sim\exp(H\,t)$, with $H=H_{I}\simeq constant$ representing the approximately constant Hubble parameter, for which the cosmological data imply the following [44]:
@@
Thus, the time–time component of the Ricci tensor equals approximately
@@
In the presence of primordial GW perturbations, the gravitational CS (gCS) anomaly term in (102) can condense at the end of the axion-dominated (preinflationary) stiff era [2]. This can drive the Universe to an inflationary phase, characterized by approximately constant gCS-anomaly condensate and Hubble parameter $H$, as well as the linear axion $b$ potential in (102), breaking the axion shift symmetry. This leads to a metastable inflation [2, 7, 8], due to the existence of imaginary parts in the condensate [8], which leads to a finite-lifetime inflationary era, and eventual exit from it. However, although from a purely dynamical system point of view [7, 45], it appears that it is the $b$-field that drives inflation, this is deceptive. The condensate itself is a non-linear function of the (slowly varying) Hubble parameter $H$, depending on the fourth (and even higher)-order powers of it. In this sense, it is such non-linearities of the gravitational sector that drive an inflation of the RVM type [15, 14, 13, 10, 9] [^11] For completeness, we mention that the inflationary exit in such cosmologies, which in general do not require external inflaton fields, is characterized by prolonged reheating periods [46, 47], and perhaps early-matter dominance epochs, preceeding standard radiation. This may affect primordial black hole populations and thus gravitational wave profiles during the early radiation era, with potentially detectable signatures in future intrerferometers [48, 49]..
To estimate the gCS condensation, the authors of [7], whose approach we follow here, considered tensor perturbations in the context of the effective gravitational Lagrangian (102), with $\lambda=0$. They quantize them within an effective weak gravitational field theory action [50, 51], following a canonical quantization approach to estimate the condensate, to a leading (linear) order in a perturbative expansion in the small parameter $\dot{b}H^{2}\ll 1\,.$
In the presence of the $\lambda$-dependent correction terms (IV) in (102), higher-order corrections proportional to $(\dot{b})^{2}$ are expected in the condensate, which will be shown below to be subleading. For the $\lambda=0$ case, as argued in [2], and also confirmed by the dynamical system analysis of the StRVM inflation of [7], as well as reviewed below, when there is a gravitational anomaly condensate, a spontaneous-Lorentz-violating solution of the $b$-axion equations of motion exists in which $\langle\mathcal{K}^{0}\rangle\simeq{\rm constant}$. We parametrize such constant quantity by [2, 7]
@@
during inflation, where $\varepsilon$ (not to be confused with $\epsilon$ in (IV)) is a numerical coefficient, which in the study of [8] is found to be of order $\mathcal{O}(10^{-2}$), and the Hubble parameter $H\approx H_{I}=constant$, satisfying (121). As we shall demonstrate below, the order of magnitude of (124) will not be affected by the presence of the $\lambda$-dependent term in (102), whose contributions to gCS condensate are extremely suppressed.
To see this, we should use perturbation theory in $\dot{b}$ when evaluating the gCS condensate, following [7]. To this end, formally, we consider weak tensor (GW) perturbations to the FLRW metric, and work in the transverse–traceless (TT) gauge [50, 51], which implies that only their spatial components, $h_{ij}$, $i,j=1,2,3$, are present in the gauge-fixed effective action. The respective (3+1)-dimensional line element is of the following form:
@@
with $a(t)$ as the scale factor.
On the helicity ($L,R$) basis, the metric reads
$g_{\mu\nu}=\begin{pmatrix}-1&0&0&0\\ 0&a^{2}(t)\left(1+\frac{h_{L}(t,z)+h_{R}(t,z)}{\sqrt{2}}\right)&ia^{2}(t)\left(\frac{h_{L}(t,z)-h_{R}(t,z)}{\sqrt{2}}\right)&0\\ 0&ia^{2}(t)\left(\frac{h_{L}(t,z)-h_{R}(t,z)}{\sqrt{2}}\right)&a^{2}(t)\left(1-\frac{h_{L}(t,z)+h_{R}(t,z)}{\sqrt{2}}\right)&0\\ 0&0&0&a^{2}(t)\end{pmatrix}$ (126)
@@
We consider the GW perturbations $h_{ij}$ to be small, propagating along the $z$ direction for concreteness. We thus obtain the following expression for the gCS condensate in the linearized approximation and in conformal time (the derivative with respect to which are denoted by primes) [52, 53, 7]:
@@
As discussed in [7], the quantity inside the square brackets $\left[\dots\right]$ on the right-hand side is independent of the scale factor $a$ of the Universe; thus, the gravitational CS condensate (128) is inversely proportional to the fourth power of the scale factor $a$.
To estimate (128), we first need to determine solutions to the linearized GW equations in the modified effective action (102), which read:
@@
We observe from (V) and (V) that the presence of the Chern–Simons coupling causes the equations for left and right waves to be different, a phenomenon known as cosmological birefringence. On the other hand, as expected, the $\lambda$-dependent corrections (IV) do not exhibit that property. Thus, such terms alone (i.e., in the absence of the gravitational anomaly CS term, when $\epsilon=0$) yield zero contributions to (128). Of course, as we shall argue below, there are non-trivial but suppressed $\lambda$-dependent corrections if $\epsilon\neq 0$.
To estimate these corrections, we should first solve, following [7, 8], the above equations in order to determine the so-called mode functions for the two GW polarizations. Then, upon canonical quantization of $h_{L,R}$, which now become quantum operators $\widehat{h}_{L,R}$, we can estimate the vacuum expectation value of (128) with respect to an appropriate quantum vacuum state, which, as in the $\lambda=0$ case of [7, 8], is taken to be the Bunch–Davies vacuum. This procedure will lead to a non-vanishing vacuum approximately constant expectation value for the Pontryagin term (128) during the inflationary period (indicated by the suffix “II”) [2, 7, 8]:
@@
This, due to the above discussion, is expected to be proportional to $\epsilon$.
Below, we provide a rather heuristic approach for the estimation of the order of the $\lambda$-dependent-term contribution to the gCS condensate (for a more rigorous approach, see Appendix D). To this end, we first assume an approximately constant $\dot{b}$ during inflation in which $H\simeq H_{I}={\rm constant}$, with the magnitude of $H_{I}$ given by observations [44], (121). In such a case, terms involving $\ddot{b}$ in the chiral GW equations of motion (V) and (V) can be ignored. By redefining
@@
the Equations (V) and (V), after straightforward manipulations, can be approximated by
@@
As discussed in Appendix D (cf. (192)), for the StRVM cosmological model of [2, 4, 7, 8], the quantity $\lambda\kappa^{2}\dot{b}^{2}=\mathcal{O}(10^{-11})\ll 1$. On the other hand, upon taking the Fourier transform of the GW linear polarizations $h_{L,R}$, assuming propagation along the $z$ direction, we may replace $\frac{1}{a}\partial_{z}$ by $i\frac{k}{a}$, where $k$ is the magnitude of the momentum vector of the GW polarization. We then have [2, 7] $\frac{k}{a}\lesssim\mu$, where $\mu$ is the UV cutoff, identified with the string scale [19] $M_{s}$, in the context of the StRVM in (100). For phenomenological reasons, specifically in order for the inflation in the StRVM to have a lifetime in the order of 50-60 e-foldings [44], the following constraint must be in operation, as implied by a dynamical system analysis of the StRVM condensate-induced inflation [7]:
@@
This is also in agreement with the study in [2].
On account of (100), (135), (124), and (121), we then obtain that the dimensionless terms $\epsilon(\frac{k}{a})\kappa^{2}\dot{b}\lesssim 8.3\times 10^{-7}$. Thus, the $\lambda$-dependent terms on the right-hand side of the wave in Equations (V) and (V) may be ignored. In this approximation, the  resulting equations, which are expressed in terms of $a^{\lambda}(t)$ ((132)), become equivalent to the corresponding wave equations of [7]. Following, then, this analysis, we may estimate the corresponding parity-violating condensate (128) in our modified gravitational theory (102) as
@@
where we keep first-order terms in a series expansion in powers of $\lambda$, and we take into account the aforementioned property of the gravitational CS anomaly condensate to depend on the inverse fourth power of the (rescaled) $a^{\lambda}$ ((132)). From (136), we therefore observe that the $\lambda$-dependent corrections will be suppressed compared to the $\lambda=0$ results of [2, 7, 8], and thus, the corresponding results regarding the inflationary era remain valid to an excellent approximation.
We now remark that in the $\lambda=0$ case, which constitutes the zeroth order correction in the $\lambda$ (i.e., Regge-slope $\alpha^{\prime}$ expansion in the string effective action [19]), the argumentation of [6], which was confirmed by the detailed analysis of [8], indicates that the gCS condensate (131) is proportional to the proper number density of GW sources $\mathcal{N}_{I}$ during the RVM inflationary era. In refs. [7, 8] the Chern–Simons condensate was estimated within a weak quantum gravity path integral formalism about a FLRW background.
In view of our result in (136), then, in our modified effective action (101), the corresponding gCS anomaly condensate will be given by
@@
to leading order in the slow-roll parameter $\alpha^{\prime}\,\kappa\,H\,{\dot{\overline{b}}}\,\ll\,1$ of the double-perturbative expansion, in powers of $\alpha^{\prime}$ and $\dot{b}$, where $H$ is the approximately constant Hubble parameter during inflation, satisfying the observational bound (121).
As already mentioned, the validity of (110) is understood. Thus, from (137), one observes that the effects of the presence of the $\lambda$-dependent terms in the modified gravitational action (102), as compared to the $\lambda=0$ case of [2, 4, 7], represent merely a very mild screening (reduction) of the strength of the effect of the GW source terms $\mathcal{N}_{I}^{\prime}\simeq\mathcal{N}_{I}(1-2\lambda\,\kappa^{2}\,\dot{b}^{2})$.
The quantity $\mu$ in (137) is a UV cutoff of the GW modes, and  the $\dots$ denote corrections implied by the last term of the integrand of (102), which we have just seen are subleading.
The existence of this condensate means that at a quantum level, the gravitational Chern–Simons term in the effective action can be expanded about the condensate as
@@
where the $\int d^{4}x\sqrt{-g}:b\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}:$ represent quantum fluctuations, for which: $\langle:\dots:\rangle=0$. As explained in [7], we can then write a “re-classicalized” effective action that has the form
$S_{E}=\int\bigg[\frac{1}{2\kappa^{2}}R-\frac{1}{2}\partial_{\mu}b\partial^{\mu}b\,+\,\frac{\epsilon}{8}b\langle R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\rangle+\frac{\epsilon}{8}:b\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}:-\frac{\lambda}{2}\partial_{\mu}b\partial_{\nu}bR^{\mu\nu}\bigg]\sqrt{-g}\,d^{4}x.$ (139)
@@
$\frac{1}{2\kappa^{2}}G_{\mu\nu}+\frac{\epsilon}{16}\langle R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\rangle g_{\mu\nu}=\frac{\epsilon}{2}\mathcal{C}_{\mu\nu}+\frac{\lambda}{2}\Theta_{\mu\nu}+\frac{1}{2}{T_{\rm kin}^{b}}_{\mu\nu},$ (140)
while the scalar equation of motion reads
@@
Upon expanding the effective action (138) about the condensate, the following is implied in general:
@@
where, as already mentioned, the $\dots:$ denote quantum fluctuations, characterized by a zero vacuum expectation value, $\langle:\dots:\rangle=0$ [^12] In the case of the gravitational CS term, the  quantum fluctuations do not contribute to the equations of motion of graviton in the FLRW background..
A constant condensate $\mathcal{K}^{0}$ (which spontaneously violates Lorentz symmetry [2] if we view the condensate as a vacuum expectation value of the topological current $\langle\mathcal{K}^{0}\rangle=\rm constant\neq 0$) is consistent with inflation, characterized by $H=\dot{a}/a={\rm constant}$ and $\ddot{a}/a={\rm constant}$, which, on account of (123), would also imply $\dot{b}={\rm constant}$, as in the case of [2]. The reader should notice that (124) implies a $H^{4}$ scaling of the Chern–Simons condensate (137) during inflation.
Upon concentrating on the condensate (which behaves classically) and ignoring the fluctuations, we obtain the following from (114) and (113):
$\frac{d}{dt}\langle\mathcal{K}^{0}(t)\rangle+3H\,\langle\mathcal{K}^{0}\rangle$ $=\langle R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\rangle^{\lambda}_{I}\,$